Question: suppose that the probability that a student will pass a math test is four-fifths, and that the students taking the test are all independent of one another. what is the probability that at least one student will pass the test.
Answer: .99840

Answer by dfrazzetto(283)


The p(at least one student passing) = 1 - p(all students fail)
assuming 4 students like last time:
p(all students fail) = (1/5)^4 = .0016
p(at least one student passing) = 1 - .0016 = 0.99840

D